# Wave pulses with unusual asymptotical behavior at infinity

### v2

@@ V Discussion and study of physical feasibility
Contrary to the case of $z-$dependence with fixed time instants, in the case of $ct-$dependence with fixed locations $z$ the tops of the curves of the imaginary parts are shifted to the right (if $t>0$) from the point where $ct$ equals to a fixed value of $z$. That is why in order for the dashed curve to touch the tops of the peaks, in the expression of $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$ the value of $\Delta$ was set to $\Delta=-a_{1}/4$. If one sets $\Delta=0$, the dashed curve exactly coincides with the dotted one. The latter presents an asymptotical behavior of $ct\cdot f\left(\rho=0,z=ct+\Delta,t\right)$ according to the following expression which is directly extracted from the series expansion of the expression in Eq. (13).
@@
Not only the total energy of the scalar field $f\left(\rho,z,t\right)$ is finite, but also the total energy of an electromagnetic field which can be proved by a relation [26] between the square integrability of the spectrum of a splash pulse and the total energy.
@@
×$\nabla\times\vec{F}-\frac{i}{c}\frac{\partial}{\partial t}\vec{F}=0,\quad\nabla\cdot\vec{F}=0.$ (31)
@@
The following results were obtained : for all possible choices of $\vec{m}$, excluding $(0,0,1)$, the $x$-components and $y$-components of electric and magnetic fields both have abnormal asymptotics when $z\rightarrow\infty$ while the $z$-components do not. Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eqs. (14), (15), (28), and (29). In other words, when propagating in the positive direction of the axis $z$ they decay as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$. If instead of the power $3/4$ another value between $1/2$ and $1$ is used in the definition of the splash pulse, the law of the decay changes correspondingly.
Fig. 7 illustrating the behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig. 3, the pulse’s maximum has become more salient, and the solid angle formed by a circle with radius equal to HWHM of the pulse is more than six times smaller than that of the scalar pulse resulting in $\Omega\approx 0.0003\,sr$ if $ct=10000$ (see Section III). This is due to the circumstance that according to Eq. (30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives which take their highest values in the region of the pulse maximum. It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to terms with abnormal asymptotics.
@@
We checked that when calculated $\vec{v}$ and $p$ to first order, indeed they obey Eqs. (33) and (34), i.e., they are physical observables.
The fact that the fractional splash pulse is not unidirectional does not imply that its physical feasibility is impossible. Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$. A practical approach to address this issue is to generate unidirectional finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
@@ VI Conclusion
The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics. It is remarkable that there are several solutions that decay with propagation distance more slowly than the common inverse proportional law, yet still possess finite energy. Especially interesting is the family of fractional splash pulses. Indeed, it can be easily verified that if one changes in Eq. (13) the power $3/4$ with any other value between $1/2$ and 1, the pulse has a slower than normal decay and, at the same time, its wave function is square-integrable and has finite total energy. These characteristics are passed on to physical electromagnetic as well as to acoustic fields derivable from the scalar wavefunctions of the fractional splash pulses. Hence, the fields with abnormally slow decay in the far zone can in principle be implemented as near-cycle pulses of radiofrequency, optical, or acoustical waves.
Figure 1: The real and imaginary parts of $U\left(\rho,z,t\right)$ at the instants $ct=0$ and $ct=4$ ($c\equiv 1$). The axis $x$ represents any axis transverse to the propagation axis $z$. The sign of the imaginary part has been reversed. The pulse width parameters are chosen as follows: $ct_{s}=0.3$ and $z_{s}=0.1$.
Figure 2: ”Top view” of the 3D plot of the real part of $U\left(\rho,z,t\right)$ at four instants $t$. See caption of Fig. 1.
Figure 4: Radial dependence of the imaginary parts of the expanding and converging splash waves defined in Eq. (21) and Eq. (22) at instants $ct=30$, $ct=170$, and $ct=-100$. The first two curves have been multiplied by $-R$, the third curve by $-50R$, and the last curve by $-R$. The pulsewidth parameter $ct_{s}=2$.
Figure 5: Radial dependencies of the real parts of the waves defined in Eq. (24), Eq. (25), and Eq. (23) at instants $ct=30$ and $ct=160$. All curves have been multiplied by $R$. The pulse width parameter is $ct_{s}=2$. For comparison, a logarithmically diverging dependence is shown by the dotted curve, whereas the factor $1/2$ has been inserted in accordance with the presence of the same factor in Eqs. (24) and (25)
Figure 6: Solid curves—time dependencies of the imaginary parts (multiplied by $ct$) of the fractional splash pulse $f\left(\rho=0,z,t\right)$ at points $z=30$, $z=90$, and $z=150$; dashed curve—an analogue of the expression under $lim$ in Eq. (2), i.e., $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$. For comparison, the dotted curve shows a diverging dependence $\sin(3\pi/8)\sqrt[4]{2ct}/2a_{1}$. The pulse width parameters are $a_{1}=3$ and $a_{2}=2$. $\Delta=-3/4$. The two last curves are shown only for the region $ct>0$.
Figure 7: The $x$-component of the real electric field $E_{x}\left(x,y=0,z,t\right)$ at $ct=10$ calculated from the Hertz vector (1,1,0)$f\left(\rho,z,t\right)$. The pulse width parameters are $a_{1}=1$ and $a_{2}=2$. Cf., also, with Fig. 3.

### v3

@@ 5 Discussion and study of physical feasibility
Figure 6: Solid curves—time dependencies of the imaginary parts (multiplied by $ct$) of the fractional splash pulse $f\left(\rho=0,z,t\right)$ at points $z=30$, $z=90$, and $z=150$; dashed curve—an analogue of the expression under limlim in Eq. (3), i.e., $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$. For comparison, the dotted curve shows a diverging dependence $\sin(3\pi/8)\sqrt[4]{2ct}/2a_{1}$. The pulse width parameters are $a_{1}=3$ and $a_{2}=2$. $\Delta=-3/4$. The two last curves are shown only for the region $ct>0$.
@@
Contrary to the case of $z$-dependence with fixed time instants, in the case of ctct-dependence with fixed locations $z$ the tops of the curves of the imaginary parts are shifted to the right (if $t>0$) from the point where $ct$ equals to a fixed value of $z$. That is why in order for the dashed curve to touch the tops of the peaks, in the expression of $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$ the value of $\Delta$ was set to $\Delta=-a_{1}/4$. If one sets $\Delta=0$, the dashed curve exactly coincides with the dotted one. The latter presents an asymptotical behavior of $ct\cdot f\left(\rho=0,z=ct+\Delta,t\right)$ according to the following expression which is directly extracted from the series expansion of the expression in Eq. (14).
@@
Not only the total energy of the scalar field $f\left(\rho,z,t\right)$ is finite, but also the total energy of an electromagnetic field which we proved by numerical double integration and analytical calculations with the help of spectral representation of fractional splash pulses (see Appendix).
@@
$\nabla\times\vec{F}-\frac{i}{c}\frac{\partial}{\partial t}\vec{F}=0,\quad\nabla\cdot\vec{F}=0.$ (31)
@@
The following results were obtained: for all possible choices of $\vec{m}$, excluding $(0,0,1)$, the $x$-components and $y$-components of electric and magnetic fields both have abnormal asymptotics when $z\rightarrow\infty$ while the $z$-components do not. Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eq.s (15), (16), (28), and (29). In other words, when propagating in the positive direction of the axis $z$ they decay as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$. If instead of the power $3/4$ another value between $1/2$ and $1$ is used in the definition of the splash pulse, the law of the decay changes correspondingly.
[…]
Fig. 7 illustrating the behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig. 3, the pulse’s maximum has become more salient, and the solid angle formed by a circle with radius equal to HWHM of the pulse is more than six times smaller than that of the scalar pulse resulting in $\Omega\approx 0.0003\,sr$ if $ct=10000$ (see Section 3). This is due to the circumstance that according to Eq. (30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives which take their highest values in the region of the pulse maximum. It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to terms with abnormal asymptotics.
@@
We checked that when $\vec{v}$ and $p$ are calculated to first order, indeed they obey Eq.s (33) and (34), i.e., they are physical observables.
The fact that the fractional splash pulse is not unidirectional does not imply that its physical feasibility is impossible. Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$. A practical approach to generate a unidirectional finite-aperture approximation of the pulse is using the Huygens approach [5], or following a method similar to what was done in the experimental realization of the focus wave mode [30, 31]. As pointed out in [5], the resulting finite-energy unidirectional fractional splash mode pulse may not retain the abnormal decay along the z-direction as $z\rightarrow\infty$. There will be, however, an intermediate zone, roughly up to the Rayleigh distance , where the pulse will be characterized by slow decay before it assumes the usual $1/z$ behavior in the far field. If the aperture radius $r_{a}$ is equal to the focused radius of the pulse at $z=0$, then the Raylegh diffraction limit is given by $Z_{R}=(\omega_{max}r_{a}^{2})/(2c)$, where $\omega_{max}=4c/a_{1}$ is the maximum effective angular frequency in the case of the fractional pulse. It is clear, then, that for a small value of the free parameter $a_{1}$ and a large aperture the region of abnormal (slow) decay can be very large. Generally, the larger the aperture, the longer the distance over which abnormally slow decay occurs. In a sense, this is analogous to an apertured Bessel beam, where the distance of non-diffracting propagation is not infinite but still increases with the size of the aperture. We can conclude that wave regions situated far from the propagation axis at $z=0$ contribute to the phenomenon of abnormally slow decay.
@@ 6 Conclusion
The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics, optics, electromagnetics and acoustics. It is remarkable that there are several solutions that decay with propagation distance more slowly than the common inverse proportional law, yet still possess finite energy. Especially interesting is the family of fractional splash pulses. Indeed, it can be easily verified that if one changes in Eq. (14) the power $3/4$ with any other value between $1/2$ and 1, the pulse has a slower than normal decay and, at the same time, its wave function is square-integrable and has finite total energy. These characteristics are passed on to physical electromagnetic as well as to acoustic fields derivable from the scalar wavefunctions of the fractional splash pulses. Hence, the fields with abnormally slow decay in the far zone can in principle be implemented as near-cycle pulses of radiofrequency, optical, or acoustical waves.
The solutions presented here complement the extensive body of work on localized, nondiffracting, spatiotemporal, and autofocusing waves, which have already found applications in fields such as particle manipulation, micromachining, nonlinear spectroscopy, data communication and storage, microscopy, etc., up to medical diagnostics and therapy. Since, e.g., the fractional splash pulses share some essential properties with the listed waves and additionally exhibit exceptionally slow intensity decay and angular broadening during propagation, they have promising prospects for applications in the same fields.
To summarize, we hope that abnormally decaying pulses as an emerging subfield in the study and applications of the localized space-time wave packets, which so far is represented to our best knowledge only by Refs. [3, 12, 5, 13, 14] and the present paper, will attract growing interest.
APPENDIX.
Proof of square integrability and finiteness of total energy of fractional splash pulses
Here we present a thorough proof of the square integrability of fractional splash pulses and finiteness of their total energy. We rely on the approach developed in Refs.[7, 32], but our derivation of the key integral relation is simpler.
A general splash pulse is given by a spectral superposition
$f_{G}(\rho,z,t)=\int_{0}^{\infty}G_{FWM}(\rho,z,t)F_{\nu}(k)dk\;,$ (35)
where
$G_{FWM}(\rho,z,t)=\frac{\exp\left\{-k\rho^{2}/\left(a_{1}+i\tau\right)+ik\left(z+ct\right)\right\}}{a_{1}+i\tau}$ (36)
is the wave function of the focus wave mode, $\tau=z-ct$, and $a_{1}$ is a positive constant with the dimension of lenght. $F_{\nu}(k)$ is a spectrum of the form
$F_{\nu}(k)=\frac{1}{\Gamma\left(\nu+1\right)}k^{\nu}\exp\left(-a_{2}k\right);\;a_{2},k>0~,$ (37)
which if $\nu=-1/4$ gives the 3/4 fractional splash pulse in Eq. (14).
Following the approach of Refs.[7, 32], the square integrability of the function $f_{G}(\rho,z,t)$ is determined by finiteness of the right-hand side of the equality
$\int\limits_{0}^{\infty}d\rho\rho\int\limits_{-\infty}^{\infty}dz~\left|f_{G}(\rho,z,t)\right|^{2}=\pi\int\limits_{0}^{\infty}dk~e^{2a_{1}k}E_{1}\left(2a_{1}k\right)\left|F_{\nu}(k)\right|^{2},$ (38)
where $E_{1}\left(x\right)$ is the exponential integral function of the first order. For brevity we have omitted here the factor $2\pi$ which comes from integration over the azimuthal angle. So our task is to prove Eq. (38) and then to study whether the right-hand side is finite or diverges depending on the value of $\nu$. Inserting Eq.s (35)-(37) into left-hand side of Eq. (38), the integration over $\rho$ with setting $t=0$ (total energy does not depend on time), leaves the integration over $z$ into the form
$I=\frac{1}{2}\int\limits_{-\infty}^{\infty}dz~\frac{\exp\left[-i\left(k_{1}-k_{2}\right)z\right]}{a_{1}\left(k_{1}+k_{2}\right)+i\left(k_{1}-k_{2}\right)z}~.$ (39)
We evaluate this key integral by the following transformations. First for the sake of brevity we denote $\Delta k\equiv k_{1}-k_{2}$ and $\Lambda\equiv a_{1}\left(k_{1}+k_{2}\right)$. One can verify that $I=0$ unless $\Delta k=0$. But if we divide the integral into two parts, we obtain
$I_{1}$ $\equiv\frac{1}{2}\int\limits_{-\infty}^{0}dz~\frac{\exp\left[-i\Delta kz\right]}{\Lambda+i\Delta kz}=\frac{i}{2}\frac{\operatorname{E}_{1}\left(\Lambda\right)e^{\Lambda}}{\Delta k},$ (40)
$I_{2}$ $\equiv\frac{1}{2}\int\limits_{0}^{\infty}dz~\frac{\exp\left[-i\Delta kz\right]}{\Lambda+i\Delta kz}=-\frac{i}{2}\frac{\operatorname{E}_{1}\left(\Lambda\right)e^{\Lambda}}{\Delta k}.$ (41)
These equalities can be found by packages of scientific calculations, or proved by changes of variables of integration. For example, Eq. (41) can be proved by the chain of the following changes of variables: $z\rightarrow\Lambda t,$ $t\rightarrow-it,$ $i(1+\Delta kt)\rightarrow y,$ $y\rightarrow it$, which results in
$\frac{i}{2}\frac{e^{\Lambda}}{\Delta k}\int\limits_{1}^{\infty}\frac{\exp\left(-\Lambda t\right)}{t}=-\frac{i}{2}\frac{\operatorname{Ei}_{1}\left(\Lambda\right)e^{\Lambda}}{\Delta k},~Q.E.D.$ (42)
Here, the definition of the exponential integral function $E_{1}\left(x\right)$ has been used. Eq. (40) is proved in the same way.
Now we apply the relation
$\lim_{\varepsilon\rightarrow 0}\frac{1}{x\mp i\varepsilon}=\pm i\pi\delta(x)+P\frac{1}{x}$ (43)
known for distributions, to the right-hand sides of Eq.s (40) and (41). Common notations are used here: $\delta(x)$ is the Dirac delta and $P$ means the principal value. Applying Eq. (42) with the opposite signs to Eq.s (40) and (41) and summing them up, the principal values cancel out. Finally, returning to the initial designations we obtain
$I=I_{1}+I_{2}=\pi\delta(k_{1}-k_{2})\operatorname{E}_{1}\left[a_{1}\left(k_{1}+k_{2}\right)\right]e^{a_{1}\left(k_{1}+k_{2}\right)}~.$ (44)
From insertion Eq.s (35)-(37) into left-hand side of Eq. (38) there remains only the integration over $k_{1}$ and $k_{2}$, one of which can be trivially carried out thanks to the Dirac delta in the integrand. Denoting the other spectral variable by $k$, we obtain, finally, the right-hand side of Eq. (38). In order to study its finiteness, we make use of the upper bound $\exp(x)\operatorname{E}_{1}(x)<\ln(1+1/x)$ (Eq. 5.1.20 in Ref. [33]). Hence, with the help of Eq. (37), a splash pulse of index $\nu$ is square integrable if the integral
$B_{\nu}(a_{1},a_{2})\equiv\int\limits_{0}^{\infty}dk~\ln(1+1/2a_{1}k)\frac{\pi}{\Gamma^{2}\left(\nu+1\right)}k^{2\nu}\exp\left(-2a_{2}k\right)$ (45)
is finite. With values $1/2<\nu\leq 0$ for the integral in Eq. (45) closed-form expressions can be found, which are more or less complicated combinations of elementary and special functions that have finite real values if $a_{1},a_{2}$ are positive parameters. If $v\leq-1/2,B_{\nu}(a_{1},a_{2})$ diverges, i.e., the 1/2-fractional pulse is not square integrable, although it exhibits abnormally slow decay with distance.
Along these lines, also, finiteness of the scalar total energy of the fractional pulses with $1/2<\nu<0$ and of electromagnetic and acoustic fields derived from them can be established. However, since these energies contain spatial and temporal derivatives of the wave function $f_{G}(\rho,z,t),$ they are definitely finite without the need to calculate them, because derivatives mean multiplication by $k$ in the spectral domain, and thus remove the singularity at $k=0$ in Eq. (45) making the finiteness of it obvious.
Square integrability and finiteness of the total energies, as well as the independence of them on time, have also been proved by direct numerical 2D integrations. For example, numerical integration in the left-hand side of Eq. (38), multiplied by $2\pi$ (due to the angular integral), in the case of a particular 3/4-fractional pulse ($\nu=-1/4,~a_{1}=1,~a_{2}=2$) yields a numerical result 29.041 independently from time instants $t=0,10,100$. For this case, the Mathematica package allows to find a closed-form expression for the right-hand side of Eq. (38), whose numerical value turns out to be equal to 29.041 as well. For the same case, the numerical value of the complicated closed-form expression of the integral in Eq. (45), multiplied also by $2\pi$, turns out to be 33.206, i.e., slightly larger as it should be for an upper bound.
For the sake of completeness, we checked the finiteness of the total energy of the scalar field $f\left(\rho,z,t\right)$ also analytically. For this the modulus squared in the left-hand side of Eq. (38) was replaced by the scalar energy density
$w=\frac{1}{2}\left|\frac{\partial}{\partial ct}f\left(\rho,z,t\right)\right|^{2}+\frac{1}{2}\left|\nabla f\left(\rho,z,t\right)\right|^{2}$ (46)
and the right-hand side by the same quantity expressed through the spectral representation Eqs. (35)-(36). In the case of $\nu=-1/4$ the spectral integration in the righ-hand side results in a combination of $\sinh^{-1}$ and gamma functions, the numerical value of which for specified values of parameters ($a_{1}=1,~a_{2}=2$) turns out to be equal to the result of numerical integration on the left-hand side, i.e., which proves the finite value of the total energy.
Finiteness of the total energy of the electromagnetic fields, derived from $f\left(\rho,z,t\right)$ by the Hertz vector technique, was also proved analytically with the help of Eq. (B14) in [32]. Again, this conclusion is natural because the EM field expressions contain double derivatives of $f\left(\rho,z,t\right)$.
[…]
Funding No funding was used for the presented research.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data availability No experimental data were generated or analyzed in the presented research.